OpenStudy (anonymous):

Help please? Solve by completing the square. x^2-10x+25=144

OpenStudy (zzr0ck3r):

\(x^2+bx+c=0\\(x+\frac{b}{2})^2+c=(\frac{b}{2})^2\)

OpenStudy (anonymous):

I have x^2-10x+25=144 x^2-10x+25-144=0 x^2-10x-119=0 from there on im kinda stuck to be honest. i dont even know if that first part is right

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

x^2-10x+25 factors to (x-5)(x-5) which is the same as (x-5)^2

jimthompson5910 (jim_thompson5910):

so x^2-10x+25=144 turns into (x-5)^2=144

OpenStudy (anonymous):

x= -7 or x=17